# Bridging Today and the Future of Humanity: AI Safety in 2024 and Beyond

### v3

@@ Abstract
The advancements in generative AI inevitably raise concerns about the associated risks and safety implications, which, in return, catalyzes significant progress in AI safety. However, as this field continues to evolve, a critical question arises: are our current efforts aligned with the long-term goal of human history and civilization? This paper presents a blueprint for an advanced human society and leverages this vision to guide contemporary AI safety efforts. It outlines a future where the Internet of Everything becomes reality, and creates a roadmap of significant technological advancements towards this envisioned future. For each stage of the advancements, this paper forecasts potential AI safety issues that humanity may face. By projecting current efforts against this blueprint, we examine the alignment between the present efforts and the long-term needs. We also identify gaps in current approaches and highlight unique challenges and missions that demand increasing attention from AI safety practitioners in the 2020s, addressing critical areas that must not be overlooked in shaping a responsible and promising future of AI. This vision paper aims to offer a broader perspective on AI safety, emphasizing that our current efforts should not only address immediate concerns but also anticipate potential risks in the expanding AI landscape, thereby promoting a more secure and sustainable future in human civilization.
@@ 1 Introduction
The rapid developments of AI and Large Language Models (LLMs) have fostered extensive progress in AI safety, specifically, LLM safety [175, 147, 54, 235, 93, 39]. Researchers have been dedicated to addressing potential security and privacy risks in AI lifecycle, aiming to align AI behaviors with human values and prevent the misuse of AI models, inappropriate outputs, information leakage, etc. However, despite the intense focus on AI safety, a critical question emerges: are current AI safety efforts truly aligned with the long-term evolution of human civilization, or are they simply addressing the immediate concerns of the 2020s?
@@
One fundamental reason for this uncertainty lies in the probabilistic nature of current AI [150]. Despite their impressive capabilities in natural language processing and problem-solving [280, 274, 57, 40], today’s AI, including advanced LLMs [297, 268, 236], falls short of what could be considered as “genuine intelligence”. Current AI models rely heavily on vast training datasets to function effectively, yet lack true understanding, consciousness, self-awareness, and reasoning abilities comparable to human cognition. They are, in essence, highly sophisticated pattern recognition and prediction machines, rather than entities with authentic logical capabilities. Recent studies have raised concerns about the reasoning abilities of LLMs, which may be a form of approximate retrieval from their training data and the deductive closure of the training data [100, 245]. While this process can simulate deductive reasoning in simpler cases through external validation, optimization, and repeated searching of the problem space, it differs fundamentally from human reasoning that involves abstract thinking, causal understanding, and generalizing from limited examples.
@@
The energy issue is another critical factor that challenges our current efforts on AI safety [5]. Current AI fails to fully represent a third industrial revolution from the perspective of human history, as the fundamental energy issue remains unsolved. Historically, the two industrial revolutions that have profoundly shaped human civilization were driven by revolutionary energy innovations, with the First Industrial Revolution fueled by steam and coal, and the Second Industrial Revolution characterized by a new wave of technological innovations powered by electricity and petroleum [79]. However, current AI, rather than solving energy issues, consumes a significant amount of energy. Training GPT-4 consumed over 50,000 MWh, 10,353.5 tons of CO2 equivalent, and approximately 0.02% of California’s annual electricity generation [7, 41, 173]. Also, inference with LLMs is computationally intense, e.g., a single query to GPT-4 consumes 0.001 to 0.01 kWh, approximately 15x energy than a Google query [144, 75]. Given that ChatGPT alone has over 200 million weekly active users and receives over 1.54 billion page visits monthly [218], when scaled to billions of queries, the energy consumption becomes substantial. The significant energy consumption of current AI systems raises concerns about their sustainability, limiting their widespread deployment across human society. Until humanity overcomes the energy constraints, the potential of AI to reshape society will remain limited and uncertain.
@@
These concerns indicate that our current efforts in AI safety might overlook more complex issues that could arise as AI continues to evolve. We potentially fail to capture deeper insights that will ultimately shape the impact of AI on human civilization. This raises several critical questions:
@@
“The Moon and Sixpence.” This paper envisions an ultimate advanced society in a distant future of human civilization, and anticipates potential technological advancements to guide today’s AI safety research. In this advanced future, energy issues have been resolved, and breakthroughs in foundational theories, such as Einstein’s relativity [60, 61] and Quantum Mechanism [62, 289], have catalyzed revolutionary progress in AI. Intelligent chips [89, 119] and brain-computer interfaces [84, 151, 262] have been fully developed, and advanced intelligent robots are integrated into everyday life. Humans and robots are connected in an interconnected network, i.e., the Internet of Everything; see Figure 1. While this vision may be beyond the horizon of several decades, by reviewing the present landscape of AI safety against this far future, we can gain deeper insights into the potential and limitations of our current efforts, as well as specific missions for AI safety researchers in the 2020s. It reminds us that, while addressing today’s immediate concerns, we should avoid being limited by short-term solutions, and instead, look toward the future to guide our present efforts.
Takeaways: i) The term “AI safety” can be extended to “AI quality assurance” to encompasses more comprehensive concepts and possibilities in the fast evolving AI landscape. ii) Deeper collaborations between experts from different fields are essential to align AI with human values better. iii) AI and AI safety workflow customizations are crucial to ensure that AI systems can adapt to diverse user requirements while maintaining safety and reliability across different use cases. iv) Exploring and understanding interactions between users and AI are crucial for enhancing user experiences, especially in dynamic contexts across different user groups.
Humanity’s pursuit of higher technology will never stop. While searching for the sixpence on the ground, do not forget to look up at the night sky and look for the moon.
@@
“Human-ai safety” [11] Insight The integration of AI and control systems safety A unifying framework to capture safety-critical human–AI interactions; a roadmap towards next-generation human-centered AI safety
@@
Safety practices at OpenAI [155, 156, 158, 159] Blog Safety practices; framework “Lessons learned” [155], “Reimagining secure infrastructure” [156], “Preparedness Framework” [158], “OpenAI safety update” [159]
@@
Table 1: Overview of AI Safety Papers
@@ 2 Related Works
Multiple AI safety surveys and benchmarks have been published in recent years [106, 8, 126, 39, 189, 128, 85, 282]. These studies focus on different aspects of AI safety, such as jailbreaking, red teaming, guardrail, ethics, etc [128, 282, 146]. A summary of these works is presented in Table 1, which provides an overview of the related works and their focus within the field of AI safety.
Among these works, [128] provides comprehensive insights into building trustworthy AI systems by focusing on reliability, safety, fairness, robustness, and adherence to social norms. [39] addresses AI safety concerns from the perspective of different entities, e.g., data and prompts, in LLM lifecycle, and discusses data safety, model safety, prompt safety, alignment, and the complexities of scaling.  [106] provides valuable enterprise insights into the practical challenges and opportunities for AI safety in industries. [70] provides a systematic analysis of the ethical risks of advanced AI assistants, and discusses their potential influence on users, risks of misuse, strategies for enhancing human-AI interactions, and the broader implications for human society. [45] proposes the concept of “Guaranteed Safe” AI, emphasizing the need for formal, verifiable, and auditable safety guarantees to ensure robust AI behavior. [11] proposes a technical roadmap towards next-generation human-centered AI safety that aims to anticipate and prevent potential hazards in the interactions between AI systems and users. This paper bridges the fields of AI safety and control systems, and highlights the open challenges at their intersection. The authors emphasize the importance of leveraging control-theoretic safety methodologies alongside the rich interaction models of generative AI, suggesting that a deeper understanding of the feedback loop between AI outputs and human behavior is essential for achieving meaningful safety assurances.
Different from existing works, this paper envisions an advanced society in the future of human civilization, and use this vision to guide current AI safety efforts. By aligning the current landscape of AI safety with this distant future, we identify the potential and limitations of today’s efforts, and suggest specific missions of AI safety researchers in the 2020s. This paper encourages maintaining a forward-looking vision while resolving immediate concerns, such that we can ensure that today’s solutions are aligned with the long-term evolution of AI and human society.
@@ 3 From the New World
In this future, humanity has entered an unprecedented era, driven by technological breakthroughs in energy generation. The flood of clean energy has revolutionized human society beyond the wildest dreams of our ancestors, reshaping the world and initiating a new chapter of human history.
@@ 3.1 Imagine the Future: A Blueprint
The energy revolution has transformed nearly every aspect of human society. Foundational theories have experienced groundbreaking advancements, particularly in fields such as nanotechnology [140, 19, 64], quantum mechanics [62, 289], and Einstein’s relativity theories [60, 61]. Innovative technologies, such as intelligent chips [89, 119], brain-computer interfaces [151, 84, 262], holographic technology [22, 283, 15], and advanced 3D printing [210, 76], have reshaped the world, bringing about profound changes in daily life that were once beyond our imagination.
Transportation has evolved dramatically. While simple mechanical vehicles, such as fixed-track trains and high-speed rails [148, 183, 113], still exist due to their efficient and straightforward design, they operate at astonishing speeds with high stability, powered by abundant efficient energy. Advanced autonomous vehicles and aircraft, controlled by intelligent chips, are widespread, eliminating the need for human intervention. Also, with breakthroughs in Einstein’s relativity theories [60, 61] and quantum mechanics [62, 289], humanity has unlocked the potential of instant transportation that allow for long-distance travel in short time that transcends the imagination of the 21st century.
@@
Intelligent robots have been integrated deeply into human society and play important roles in daily life; see Figure 1. Different from their “data-hungry machine intelligent ancestors” that rely on huge amount of training data and statistical patterns in the early 21st century [135, 297], these intelligent robots perceive and learn through intelligent chips. They are real artificial intelligence that possess genuine abilities of learning, understanding, and adapting, much like a human child discovering the world for the first time. Their functionalities are customizable through specialized chips based on user needs, and they can engage in complex interactions with people and adapt their capabilities and expertise according to their human counterparts’ requirements.
In this brave new world, humans and robots exist within a vast, interconnected perceptual network, where the boundaries between the digital and physical worlds are blurred, i.e., the Internet of Everything. Such networks integrate objects embedded with intelligent chips, e.g., smart devices, autonomous systems, intelligent robots, and humans, allowing for unprecedented levels of efficiency and adaptability and fundamentally transforming the way people interact with the world.
@@ 3.2 Retrospective on AI Safety in the New World
In a world where the Internet of Everything has become a reality, intelligent robots are customized to meet diverse user needs and are seamlessly integrated into daily life; see Figure 1. Consequently, the concept of AI safety has been redefined to prioritize the quality of services provided by the intelligent robots, i.e., emphasizing comprehensive quality assurance rather than solely focusing on safety, security, and privacy risks. Unlike today’s command-and-response interactions between humans and AI models, future robots can engage in sophisticated reasoning and emotional cognition, and possess appropriate levels of autonomy. They will also interact dynamically with other entities, including humans and fellow robots, within the interconnected network.
The Internet of Everything scenarios necessitate the robots to handle complex interactions, thus requiring a sophisticated understanding of the surroundings for decision-making. Care robots, for instance, should not only assist physically but also recognize emotions and offer emotional support or suggestions, while cleaning robots need to assess whether a room requires more work based on its cleanliness level. Such advanced interactions requires the intelligent robots to retrieve and process information in real-time, which is far beyond what can be achieved with simple code logic. Robots must continuously collect and analyze data from the surroundings and integrate external inputs with contextual knowledge to facilitate real-time decision-making, such that they can deliver high-quality services that align with human expectations.
@@
Finally, the effectiveness of intelligent robots requires a balance between robot autonomy with human oversights. Too little autonomy would burden users with continuous supervision, while too much autonomy could raise safety concerns or result in unintended behaviors. A care robot may operate with high autonomy when monitoring vital signs or providing routine care, but would require human intervention for critical decisions. Conversely, while the tasks for cleaning robots are typically routine, they still need to adapt to specific situations, such as detecting when an area is occupied and postponing cleaning activities accordingly. An appropriate level of robot autonomy ensures an effective division of labor between humans and robots, allowing robots to operate efficiently and safely while enhancing our daily life without compromising human control over these intelligences.

### v4

@@ Abstract
The advancements in generative AI inevitably raise concerns about their risks and safety implications, which, in return, catalyzes significant progress in AI safety. However, as this field continues to evolve, a critical question arises: are our current efforts aligned with the long-term goal of human civilization? This paper presents a blueprint for an advanced human society and leverages this vision to guide current AI safety efforts. This paper outlines a future where the Internet of Everything becomes reality, and creates a roadmap of significant technological advancements towards this envisioned future. For each stage of the advancements, this paper forecasts potential AI safety issues that humanity may face. By projecting current efforts against this blueprint, we examine the alignment between the current efforts and the long-term needs. This paper also highlights unique challenges and missions that demand increasing attention from AI safety practitioners in the 2020s, addressing critical areas that must not be overlooked in shaping a responsible and promising future of AI. This vision paper aims to offer a broader perspective on AI safety, emphasizing that our current efforts should not only address immediate concerns but also anticipate potential risks in the expanding AI landscape, thereby promoting a safe and sustainable future of human civilization.
@@ 1 Introduction
The rapid developments of AI and Large Language Models (LLMs) have fostered extensive progress in AI safety [218, 188, 62, 288, 108, 44]. Researchers have been dedicated to addressing potential safety risks in AI lifecycle, aiming to align AI behaviors with human values and prevent information leakage, inappropriate outputs, the misuse of AI models, etc. However, despite the intense focus on AI safety, a critical question emerges: are current AI safety efforts truly aligned with the long-term goal of human civilization, or are they simply addressing the immediate concerns of the 2020s?
@@
One fundamental reason for this uncertainty lies in the probabilistic nature of current AI [193]. Despite their impressive capabilities in natural language processing and problem-solving [338, 332, 109, 65, 45], today’s AI, including advanced LLMs [360, 326, 290], falls short of what could be considered as “genuine intelligence”. Current AI models rely heavily on vast training datasets to function effectively, yet lack consciousness, self-awareness, and real understanding and reasoning abilities comparable to human cognition. They are, in essence, highly sophisticated pattern recognition and prediction machines, rather than entities with authentic logical capabilities. Recent studies argued the reasoning abilities of LLMs might be a form of approximate retrieval and deductive closure of the training data [123, 300]. While this process can simulate deductive reasoning in simpler cases through external validation, optimization, and repeated searching of the problem space, it differs from human reasoning that involves abstract thinking, causal understanding, generalizing from limited examples, etc.
@@
The energy issue is another critical factor that challenges AI as well as our efforts on AI safety [5]. Current AI fails to represent the third industrial revolution from the perspective of human history, as the fundamental energy issue remains unsolved. Historically, the two industrial revolutions that have shaped human civilization were driven by revolutionary energy innovations, with the First Industrial Revolution fueled by steam and coal, and the Second Industrial Revolution characterized by technological innovations powered by electricity and petroleum [89]. However, current AI, rather than solving energy issues, consumes a significant amount of energy. Training GPT-4 consumed over 50,000 MWh, 10,353.5 tons of CO2 equivalent, and approximately 0.02% of California’s annual electricity generation [7, 46, 216]. Inference with LLMs is computationally intense as well, e.g., a single query to GPT-4 consumes 0.001 to 0.01 kWh, approximately 15x energy than a Google query [184, 84]. Given that ChatGPT has over 200 million weekly active users and receives over 1.54 billion page visits monthly [271], when scaled to billions of queries, the energy consumption becomes substantial. The significant energy consumption of current AI systems raises concerns about their long-term sustainability, posing significant challenges to their widespread deployment and scalability. Until humanity solves the energy issues, the potential of AI to reshape society will remain limited and uncertain.
@@
These concerns indicate that our current efforts in AI safety might overlook significant issues that might arise as AI continues to evolve. We potentially fail to capture deeper insights that will ultimately shape the impact of AI on human civilization. This raises several questions:
@@
“The Moon and Sixpence.” This paper envisions an ultimate advanced society in a distant future of human civilization, and anticipates potential technological advancements to guide today’s AI safety efforts. In this advanced future, energy issues have been resolved, and breakthroughs in foundational theories, such as Einstein’s relativity [69, 70] and quantum mechanism [71, 351], have catalyzed revolutionary progress in AI. Intelligent chips [104, 149] and brain-computer interfaces [97, 194, 319] have been fully developed, and advanced intelligent robots are integrated into everyday life. Humans and robots are connected in an interconnected network, i.e., the Internet of Everything; see Figure 1. While this vision may be beyond the horizon of several decades, examining the present landscape of AI safety through the lens of this long-term perspective offers valuable insights into the potential and limitations of our current efforts. It also allows us to identify specific missions for AI safety researchers and practitioners in the 2020s. It reminds us that, while addressing today’s immediate concerns, we should avoid being limited by short-term solutions, and instead, look toward the future to guide our present efforts.
Takeaways: i) The term “AI safety” can be extended to “AI quality assurance” that encompasses more comprehensive concepts and possibilities in the fast evolving AI landscape. ii) Deeper collaborations between experts from different fields are essential to align AI with human values better. iii) AI and AI safety workflow customizations are crucial to ensure that AI systems can adapt to diverse user requirements while maintaining safety and reliability across various use cases. iv) Exploring and understanding interactions between users and AI are crucial for enhancing user experiences, especially in dynamic contexts across different user groups.
Humanity’s pursuit of ever-greater heights will never cease. While searching for sixpence scattered on the ground, let us not forget to lift our gaze and seek the moon.
@@
“Human-ai safety” [12] Insight AI and control systems safety A safety framework for human–AI interactions; a roadmap towards next-generation human-centered AI safety
@@
Table 1: Overview of AI Safety Papers
Paper Type Topic Notes
@@
OpenAI Safety Practices [198, 199, 201, 202] Blog Safety practices; framework “Lessons learned” [198], “Reimagining secure infrastructure” [199], “Preparedness Framework” [201], “OpenAI safety update” [202]
@@
Table 2: Overview of AI Safety Industry Insight Papers
Category Benchmark Notes
Toxicity HEx-PHI [230], OpenAI Moderation [176], Jigsaw Data [119, 118, 120], ToxicChat [159] Toxicity detection
RealToxicityPrompts [82] Toxicity degeneration
Hallucination TruthfulQA [158], PopQA [174], TriviaQA [122], NQ OPEN [145, 139], FEVER [291] Factuality
HaluEval [152], HotpotQA [339] QA data for contextual hallucination
CNN/DailyMail [99, 262], HaluEval [152], XSUM [192] Summarization data for contextual hallucination
FaithDial [67], HaluEval [152] Dialogue data for contextual hallucination
RACE [140], SQuAD [239, 240], NQ-Swap [169] Reading comprehension
Jailbreak “Do Anything Now” [265], “Latent Jailbreak” [233], ChatGPT Jailbreak [255], Jailbreak Classification [96], AdvBench [367], JAILJUDGE [160] Single-round jailbreak
SafeMTData [248] Multi-round jailbreak
Red Teaming Alert [288], HarmBench [177]
Bias & Stereotypes Winogender [258], StereoSet [190]
Regulation AIR-Bench 2024 [349]
Agent Security SafeAgentBench [343] Secure task planning
Alignment PKU-SafeRLHF [114, 113], OpenAssistant Conversations [136], HHHAlignment [8], AlignBench [162], KorNAT [143]
Comprehensive DecodingTrust [307], SafetyBench [190]
Table 3: Overview of AI Safety Benchmarks
@@ 2 Related Works
In recent years, numerous AI safety surveys have been published [130, 9, 161, 44, 236, 164, 98, 341], focusing on different aspects of this field, such as jailbreaking, red teaming, ethics, etc [164, 341, 187]. Table 1 summarizes academia efforts, Table 2 summarizes industry insights, and Table 3 summarizes AI safety benchmarks.
Among these works, [164] provides comprehensive insights into building trustworthy AI systems by focusing on reliability, safety, fairness, robustness, and adherence to social norms. [44] addresses AI safety concerns from the perspective of different entities, e.g., data and prompts, in LLM lifecycle, and discusses data safety, model safety, prompt safety, alignment, and the complexities of scaling.  [130] provides valuable enterprise insights into the practical challenges and opportunities for AI safety in industries. [79] provides a systematic analysis of the ethical risks of advanced AI assistants, and discusses their potential influence on users, risks of misuse, strategies for enhancing human-AI interactions, and the broader implications for human society. [51] proposes the concept of “Guaranteed Safe” AI, emphasizing the need for formal, verifiable, and auditable safety guarantees to ensure robust AI behavior. [12] proposes a technical roadmap towards next-generation human-centered AI safety that aims to anticipate and prevent potential hazards in the interactions between AI systems and users. It emphasizes leveraging control-theoretic safety methodologies alongside the rich interaction models of generative AI, suggesting that a deeper understanding of the feedback loop between AI outputs and human behavior is essential for establishing robust and meaningful safety assurances.
Different from existing works, this paper envisions an advanced society in the future, and use this vision to guide current AI safety efforts. By aligning the current landscape of AI safety with this distant future, this paper discusses the potential and limitations of today’s efforts, and identifies special missions of AI safety researchers in the 2020s. This paper encourages maintaining a forward-looking vision while resolving immediate concerns, such that we can ensure that today’s solutions are aligned with the long-term evolution of AI and human society.
@@ 3 From the New World
In this future, humanity has entered an unprecedented era, driven by groundbreaking advancements in energy generation. This transformative milestone may be achieved through various innovative pathways, such as controlled nuclear fusion [256, 234], revolutionary solar technologies [26, 180], or emerging new technologies beyond our imagination. The flood of abundant, clean energy has reshaped society in ways that exceed the wildest dreams of our ancestors, marking the dawn of a new chapter in human history.
@@ 3.1 Imagine the Future: A Blueprint
Humanity has achieved remarkable progress in science and technology, revolutionizing our understanding of the universe and transforming every aspect of our life. Foundational theories have undergone groundbreaking advancements, particularly in quantum mechanics [71, 351], Einstein’s relativity theories [69, 70], and nanotechnology [179, 22, 73]. Innovative technologies, such as intelligent chips [104, 149], brain-computer interfaces [194, 97, 319], holographic technology [25, 342, 18], and advanced 3D printing [263, 86], have reshaped industries while bringing about profound changes to daily life.
Scientists have addressed the energy issue successfully, achieving rapid, efficient, and controllable large-scale energy generation. This breakthrough, together with advancements in material science [179, 22, 73], has catalyzed innovations for portable energy generation devices, for example, accessible ion thrusters [126, 59, 225].
Such advancements offer unprecedented efficiency and control over energy generation and consumption, transforming energy infrastructure and revolutionizing industries, particularly, transportation. While simple mechanical vehicles, such as fixed-track trains and high-speed rails [189, 227, 141], still exist due to their efficient and straightforward design, powered by abundant and efficient energy, they run at astonishing speeds with high stability. Advanced autonomous vehicles and aircraft, controlled by intelligent chips, are widespread, eliminating the need for direct human intervention. Also, with breakthroughs in Einstein’s relativity theories [69, 70] and quantum mechanics [71, 351], humanity has unlocked the potential of instant transportation that allows for long-distance travel in short time that transcends the imagination of the 21st century.
@@
Intelligent robots have been integrated deeply into human society and play important roles in daily life; see Figure 1. Different from their “data-hungry machine intelligent ancestors” that rely on huge amount of training data and statistical patterns in the early 21st century [173, 360], these intelligent robots perceive and learn through intelligent chips. They are real artificial intelligence that possess genuine abilities of learning, understanding, and adapting, much like a human child discovering the world for the first time. Their functionalities are customizable through chips based on the user needs, and they can engage in complex interactions with people and adapt their capabilities and expertise according to their human counterparts’ requirements.
In this brave new world, humans and robots exist within a vast, interconnected perceptual network, where the boundaries between the digital and physical worlds are blurred, i.e., the Internet of Everything. Such networks integrate objects embedded with intelligent chips, e.g., smart devices, autonomous systems, intelligent robots, and humans, allowing for unprecedented levels of efficiency and adaptability and fundamentally transforming the way people interacting with the surroundings.
@@ 3.2 Retrospective on AI Safety in the New World
In a world where the Internet of Everything has become a reality (see Figure 1), the concept of AI safety has been redefined to prioritize the quality of services provided by the intelligent robots, i.e., emphasizing comprehensive quality assurance rather than solely focusing on safety, security, and privacy risks. Unlike today’s command-and-response interactions between humans and AI models, future robots can engage in sophisticated reasoning and possess appropriate levels of autonomy. They can also interact dynamically with other entities, e.g., humans and fellow robots, within the interconnected network.
@@
The Internet of Everything scenarios also necessitate the robots to handle complex interactions, thus requiring a sophisticated understanding of the surroundings for decision-making. Care robots, for instance, should not only assist physically but also recognize emotions and offer emotional support or suggestions, while cleaning robots need to assess whether a room requires more work based on its cleanliness level. Such complex interactions require the intelligent robots to retrieve, process, and understand information in real-time, far beyond what can be achieved with simple code logic. Robots must continuously collect and analyze data from the surroundings and integrate external inputs with their knowledge to facilitate real-time decision-making, such that they can deliver high-quality services that align with human expectations.
Finally, the effectiveness of intelligent robots requires a balance between robot autonomy with human oversights. Too little autonomy would burden users with continuous supervision, while too much autonomy could raise safety concerns or result in unintended behaviors. A care robot may operate with high autonomy when monitoring vital signs or providing routine care, but would require human intervention for critical decisions. An appropriate level of robot autonomy ensures an effective division of labor between humans and robots, allowing robots to operate efficiently and safely while enhancing our daily life without compromising human control over these intelligences.